# Theory uncertainties of the irreducible background to VBF Higgs production

### v1

@@ I Introduction
The discovery of the Higgs boson Aad et al. (2012); Chatrchyan et al. (2012) by the ATLAS and CMS experiments at the Large Hadron Collider (LHC) has ushered in a new era in elementary particle physics. The so far excellent agreement of the boson’s properties with Standard Model expectations leaves as many questions open as it answers, and calls for a more detailed investigation of this unique particle. One of the most promising Higgs production modes for in-depth studies is weak vector boson fusion, conventionally referred to as VBF. This process has been analyzed by both ATLAS and CMS for a number of Higgs boson decay channels at a number of center-of-mass energies Aad et al. (2022); Tumasyan et al. (2022). High-statistics measurements of VBF allow for a precise determination of the couplings of the Higgs boson to and bosons, as well as improved limits in searches for possible new physics. The presence of two (or more) jets in the final state, characterized by a large rapidity separation and/or a large dijet invariant mass, can be used to enhance the VBF sub-process over the dominant production mechanism, which is gluon-gluon fusion (ggF). However, the final-state configurations in gluon fusion with two (or more) jets can still be identical to the one in VBF, turning ggF from a signal into an irreducible background. The full potential of the LHC for VBF Higgs boson measurements can therefore only be realized, once gluon fusion Higgs production with two additional jets is theoretically understood to the highest achievable precision.
@@
Matrix element plus parton shower matching Frixione and Webber (2002); Nason (2004); Frixione et al. (2007); Alioli et al. (2010); Hoche et al. (2011); Höche et al. (2012); Alwall et al. (2014); Jadach et al. (2015); Nason and Salam (2022) and merging André and Sjöstrand (1998); Catani et al. (2001); Lönnblad (2002); Mangano et al. (2002); Krauss (2002); Lavesson and Lönnblad (2008a); Höche et al. (2009); Hamilton et al. (2009); Lönnblad and Prestel (2012, 2013a); Plätzer (2013); Höche et al. (2019); Lavesson and Lönnblad (2008b); Hoeche et al. (2013); Frederix and Frixione (2012); Lönnblad and Prestel (2013b); Hoeche et al. (2014) techniques are an essential theory tool to achieve this goal, and are used in the extraction of the Higgs boson cross sections from experimental measurements. They are embedded in complete frameworks for the production of the particle-level final state provided by the general-purpose LHC event generators, Herwig, Pythia, and Sherpa Buckley et al. (2011); Campbell et al. (2024). Many other theory tools make use of the event generators to provide the implementation of resummation and/or hadronization in terms of parton showers and fragmentation models. The interplay between perturbative and non-perturbative QCD is a key aspect of Monte Carlo event generation at the LHC. Simulations typically assume a scale hierarchy, where the hardest interactions of the event are described by perturbatively computed matrix elements, while the parton shower adds softer radiation using resummed perturbation theory. At some point during the QCD evolution, a scale is reached where the transition to the non-perturbative regime must be implemented. The simulation of this hadronization process relies on physically motivated models that are tuned to experimental data from current and past experiments, notably data from LEP, HERA and the LHC. As these data sets are the same for all event generators, the different hadronization models usually predict fairly similar results, and their uncertainties tend to be of a similar size Höche et al. (2012); Alcaraz Maestre et al. (2012), although in cases where particle spectra are resolved in detail, significant differences can still occur Carli et al. (2010); Divisova et al. (2025). On the other hand, one often observes discrepancies in the simulation of perturbative QCD effects. A classical example would be if the parton shower settings involved a different form of the strong running coupling than the hard matrix elements, or if the value of the running coupling differed at the reference scale. Such mismatches are due to the attempt to “tune away” effects that can arise from missing higher-order corrections. Inconsistent tunes, however, are not predictions based on solid QCD perturbation theory, and the corresponding simulation results can no longer be relied upon.
@@
In a previous publication Buckley et al. (2021), some of us focused on a consistent way to simulate VBF Higgs boson production at both fixed-order (LO, NLO, NNLO) and fixed-order plus parton shower, finding agreement in both the central predictions and their uncertainties at the level expected of the perturbative QCD result. These calculations are in line with the observations of a range of other studies Jäger et al. (2014, 2020); Ballestrero et al. (2018); Höche et al. (2022); Bittrich et al. (2022); Chen et al. (2022a); van Beekveld and Ferrario Ravasio (2024). In the present paper, we will explore the gluon-gluon fusion background, using a variety of tools in the same common setup as required to obtain theoretically consistent results.
@@
The analyses performed here match those used by ATLAS and CMS as much as possible. We will show that the nominal setups for the different programs lead to similar predictions for the kinematic distributions crucial to the determination of the VBF Higgs signal. As a result, the uncertainties quoted for the ggF backgrounds by ATLAS and CMS may be overstated. We will compare predictions from approaches that are inclusive in the number of jets (; such as from Sherpa MEPS@NLO) to those from more exclusive approaches, where only the final 2 jet state is produced (such as Sherpa S-MC@NLO and Powheg Box + Pythia/ Powheg Box + Herwig ). We also calculate the MiNNLO prediction for gluon-fusion (ggF) Higgs boson production, in which the leading jet spectrum is computed at NLO precision, while the second jet is predicted at LO. This is the framework which is currently used by the ATLAS and CMS experiments. We compare results at the parton shower level and the hadron level and discuss similarities and differences. We disentangle uncertainty estimates as much as possible, in order to provide a better understanding of their sources. We also test the impact of the merging scheme and scale. We provide a Rivet routine that can be directly used by the experiments to confirm the correct use of each generator. Although the ATLAS and CMS determinations of the VBF Higgs cross section are not strictly cut-based, we will also discuss a cutflow analysis, indicating the results of successive phase-space restrictions that are used in VBF Higgs analyses, as well as providing comparisons of what has been termed loose and tight VBF selection cuts.
@@ II Input parameters and Computational Tools
We eliminate the two antipodal combinations and . The parton distribution functions (PDFs) are defined by the central group of PDF4LHC21_40 NNLO PDFs Ball et al. (2022). The computational tools used in this study are introduced in the following.
@@ II.1 NNLOJet
NNLOJet is a parton level event generator specializing in the computation of NNLO QCD corrections Huss et al. (2025). The antenna subtraction formalism Gehrmann-De Ridder et al. (2005a, b); Daleo et al. (2007, 2010); Gehrmann-De Ridder et al. (2012); Currie et al. (2013) is used to regulate IR divergences at each stage of the fixed-order calculations and to provide fully-differential results. For Higgs production in the gluon-fusion channel, NNLOJet provides predictions at NNLO Chen et al. (2015, 2016). This implies that for observables with at least one (), two () and three jets (), the corresponding fixed order QCD accuracy is NNLO, NLO and LO, respectively. In this study, the calculation is performed in the Higgs effective field theory (HEFT) with infinitely large top-quark mass. The relevant Standard Model (SM) top-quark mass effects were studied for at NLO with full theory approximation on the two-loop matrix elements finding comparable NLO/LO ratios between HEFT and SM for differential observables Chen et al. (2022b), justifying the treatment of top-quark mass in the rest of this study. More details on this can be found in App. A.
@@ II.2 POWHEG Box
We note that the MiNNLO-based prediction for the 2-jet final state is leading-order accurate. Next-to-leading order precision is achieved by using the simulation, which requires generation cuts or Born suppression (see Sec. IV.2 for details). While Born suppression works well at the parton level, full event generation can fail if the simulation of multiple parton interactions is allowed to generate jets at transverse momenta larger than that of the hard process. Events with low transverse momentum jets in the hard process, which have large (and in principle unbounded) weights, can then be elevated to the 2-jet final state, which will result in distributions with seemingly random statistical outlier events. We therefore decide to define the upper scale for both parton-shower and MPI evolution as the value given by the SCALUP parameter. This technique is not available in Herwig, therefore comparisons including Powheg Box + Herwig will be limited to the parton level.
We have cross-checked the Powheg Box NLO predictions at fixed order against NNLOJet, Herwig and Sherpa. Note that a sample with all three multiplicities (, and ) provided at NLO accuracy can be obtained with an extension of the MiNLO method described in Ref. Frederix and Hamilton (2016), but the corresponding code is currently not publicly available.
@@ II.3 Pythia
Pythia Sjöstrand et al. (2006); Bierlich et al. (2022) is a multi-purpose particle-level event generator with a historically strong focus on soft physics. The non-perturbative evolution from colored partons to color-singlet hadrons is modeled using its distinctive feature, the string-fragmentation model Andersson et al. (1983); Sjöstrand (1984). In the present version 8.3, Pythia implements the Vincia antenna shower Brooks et al. (2020) in addition to the default -ordered “simple shower” Sjöstrand and Skands (2005)111In versions prior to 8.316, also the Dire dipole-like parton shower Höche and Prestel (2015) was available.. For the simple shower, a dipole-recoil option is available, which replaces the separate incoherent initial-final and final-initial parton evolution by a common, coherent initial-final evolution Cabouat and Sjöstrand (2018). For each of the parton-shower algorithms, multi-parton interactions are fully interleaved with the shower evolution Sjöstrand and Skands (2005). In this study, we use Pythia 8.315.
Pythia 8.3 implements a series of processes at LO. Both the VBF Higgs-production process and the gluon-fusion Higgs production process in the heavy-top limit are available at LO accuracy. In case of the latter, the simple shower algorithm also incorporates LO matrix-element corrections for the first emission, corresponding to the configuration. For matching and merging, Pythia relies on Les Houches events Boos et al. (2001); Alioli et al. (2014) from external programs such as MadGraph5_aMC@NLO Alwall et al. (2014), Powheg Box, or Sherpa. Multi-jet merging is implemented in the CKKW-L approach Lönnblad (2002); Lönnblad and Prestel (2012). For NLO matching with the Powheg Box, dedicated PowhegHooks are available for all internal parton-shower algorithms to account for the mismatch between the evolution variable in Powheg Box and the shower evolution variable in Pythia Corke and Sjöstrand (2010). NLO matching with MadGraph5_aMC@NLO is only available with the default simple shower with (non-default) global longitudinal recoil in final-state branchings, which was found to be problematic for certain simulations and will therefore not be considered in this study.
@@ II.4 Herwig
In this article, we use the angular-ordered shower to process Les Houches events generated by the Powheg Box. In particular, we use the powhegHerwig plugin developed in Ref. Ferrario Ravasio and Oleari (2023), which is largely based on the Les Houches interface provided by the ThePEG Lonnblad (2009). Independently, Herwig’s fixed-order calculation of the cross section and distributions has been checked against NNLOJet, Powheg Box and Sherpa.
@@ II.5 Sherpa
Sherpa Gleisberg et al. (2004, 2009); Bothmann et al. (2019, 2024) is a general-purpose event generator with a strong focus on perturbative physics, in particular multi-jet merging methods. We use version 3.0 of the program for this study. The matching of fixed-order NLO corrections to the parton shower is performed in the S-MC@NLO approach Höche et al. (2012, 2013), while NLO computations are performed using Amegic Krauss et al. (2002); Gleisberg and Krauss (2008), Comix Gleisberg and Höche (2008), and MCFM Campbell and Ellis (1999); Campbell et al. (2011, 2015); Campbell and Neumann (2019); Campbell et al. (2021). We have cross-checked the virtual corrections against OpenLoops Cascioli et al. (2012); Buccioni et al. (2018, 2019) and compared the complete NLO predictions against NNLOJet, Herwig and Powheg Box. We use a modified version of the Catani-Seymour dipole shower algorithm (CSS) Schumann and Krauss (2008), which was cross-checked against the Dire parton shower Höche and Prestel (2015) (for S-MC@NLO only). The latter has been deprecated in Sherpa, and therefore we only present predictions using the default CSS algorithm. The NLO multi-jet merging is based on Gehrmann et al. (2013); Hoeche et al. (2013). Hadronization in Sherpa is provided by a cluster fragmentation model Winter et al. (2004); Chahal and Krauss (2022), and multiple parton scattering as a model of the underlying event is simulated according to Alekhin et al. (2006a, b).
@@ III Comparison to typical LHC simulations
Before performing a detailed analysis of theoretical uncertainties, we compare our new predictions to those obtained from a setup routinely employed by the experimental collaborations. This provides the context for the discussion of the true uncertainties in Sec. IV. The event generator conventionally used by ATLAS and CMS for both VBF Higgs boson production and the ggF Higgs production background was NNLOPS Hamilton et al. (2013a, 2015). It has now largely been replaced by the more modern MiNNLO generator Monni et al. (2020); Niggetiedt and Wiesemann (2024), which gives formally equivalent results while being substantially more efficient, due to the elimination of the reweighting step. Both NNLOPS and MiNNLO are usually interfaced with the Herwig and Pythia Monte-Carlo programs. The estimated uncertainty in the ggF Higgs boson background is dominated by the observed differences between the predictions of NNLOPS+Pythia and NNLOPS+Herwig, which arise from both perturbative and non-perturbative QCD effects. A key point is that these are the differences observed by using the simulations “out of the box” (albeit with tuned non-perturbative parameters from each experiment), and not differences resulting from true discrepancies in the physics models of the event generators. In this section, we will use the ATLAS and CMS setups of MiNNLO for comparison to the ones we recommend as a new baseline to reach the highest theoretical precision in jets from ggF. Note in particular that the use of NNLOPS to simulate the ggF background to VBF Higgs production implies that the final state is predicted at leading order accuracy only, which is below the current state of the art. We also stress that in order to combine predictions from a Powheg Box generator with an angular-ordered shower, one would need to include truncated showers to fill the large-angle region Nason (2004). This also applies even when Powheg Box events are generated in accordance with the NNLOPS or MiNNLO methods, where now one has to imagine the two hardest emissions having been pre-determined. Truncated showers are implemented in the internal Herwig implementation of POWHEG, but not if one showers (NLO or NNLO) events generated with external tools. One can argue that at leading color, simply vetoing radiation with a transverse momentum larger then the (softest) Powheg Box emission is sufficient to fill the large-angle region while retaining leading logarithmic accuracy (see e.g. Appendix D of Ref. Ferrario Ravasio et al. (2018)). The same reasoning applies if one considers two emissions as in the case of NNLOPS/MiNNLO. In the case of POWHEG matching, the effect of truncated showers has been studied e.g. in Ref. Hamilton et al. (2008), and found to be small, but such a study is missing in the case of two hard emissions. We give more details on the precise Monte-Carlo simulations used in the experiments in App. C.

### v2

@@ I Introduction
The discovery of the Higgs boson Aad et al. (2012); Tumasyan et al. (2022a) by the ATLAS and CMS experiments at the Large Hadron Collider (LHC) has ushered in a new era in elementary particle physics. The so far excellent agreement of the boson’s properties with Standard Model expectations leaves as many questions open as it answers, and calls for a more detailed investigation of this unique particle. One of the most promising Higgs production modes for in-depth studies is weak vector boson fusion, conventionally referred to as VBF. This process has been analyzed by both ATLAS and CMS for a number of Higgs boson decay channels at a number of center-of-mass energies Aad et al. (2021a, b, 2023a, 2022, 2025); Chatrchyan et al. (2012); Hayrapetyan et al. (2024a); Sirunyan et al. (2021a, b, 2018, 2019a); Tumasyan et al. (2022b). High-statistics measurements of VBF allow for a precise determination of the couplings of the Higgs boson to and bosons, as well as improved limits in searches for possible new physics. The presence of two (or more) jets in the final state, characterized by a large rapidity separation and a large dijet invariant mass, can be used to enhance the VBF sub-process over the dominant production mechanism, which is gluon-gluon fusion (ggF). However, the final-state configurations in gluon fusion with two (or more) jets can still be identical to the ones in VBF, turning ggF from a signal into an irreducible background. The full potential of the LHC for VBF Higgs boson measurements can therefore only be realized, once gluon fusion Higgs production with two additional jets is theoretically understood to the highest achievable precision.
@@
Matrix element plus parton shower matching Frixione and Webber (2002); Nason (2004); Frixione et al. (2007); Alioli et al. (2010); Hoche et al. (2011); Höche et al. (2012); Alwall et al. (2014); Jadach et al. (2015); Nason and Salam (2022); Höche et al. (2025a) and merging André and Sjöstrand (1998); Catani et al. (2001); Lönnblad (2002); Mangano et al. (2002); Krauss (2002); Lavesson and Lönnblad (2008a); Höche et al. (2009); Hamilton et al. (2009); Lönnblad and Prestel (2012, 2013a); Plätzer (2013); Höche et al. (2019); Lavesson and Lönnblad (2008b); Hoeche et al. (2013); Frederix and Frixione (2012); Lönnblad and Prestel (2013b); Hoeche et al. (2014) techniques are an essential theory tool to achieve this goal, and are used in the extraction of the Higgs boson cross sections from experimental measurements. They are embedded in complete frameworks for the production of the particle-level final state provided by the general-purpose LHC event generators, Herwig, Pythia, and Sherpa Buckley et al. (2011); Campbell et al. (2024). Many other theory tools make use of the event generators to provide the implementation of resummation and/or hadronization in terms of parton showers and fragmentation models. The interplay between perturbative and non-perturbative QCD is a key aspect of Monte Carlo event generation at the LHC. Simulations typically assume a scale hierarchy, where the hardest interactions of the event are described by perturbatively computed matrix elements, while the parton shower adds softer radiation using resummed perturbation theory. At some point during the QCD evolution, a scale is reached where the transition to the non-perturbative regime must be implemented. The simulation of this hadronization process relies on physically motivated models that are tuned to experimental data from current and past experiments, notably data from LEP, HERA and the LHC. As these data sets are the same for all event generators, the different hadronization models usually predict fairly similar results, and their uncertainties tend to be of a similar size Höche et al. (2012); Alcaraz Maestre et al. (2012), although in cases where particle spectra are resolved in detail, significant differences can still occur Carli et al. (2010); Divisova et al. (2025). On the other hand, one often observes discrepancies in the simulation of perturbative QCD effects. A classical example would be if the parton shower settings involved a different form of the strong running coupling than the hard matrix elements, or if the value of the running coupling differed at the reference scale. Such mismatches are due to the attempt to “tune away” effects that can arise from missing higher-order corrections. Inconsistent tunes, however, are not predictions based on solid QCD perturbation theory, and the corresponding simulation results can no longer be relied upon.
@@
In a previous publication Buckley et al. (2021), some of us focused on a consistent way to simulate VBF Higgs boson production at both fixed-order (LO, NLO, NNLO) and fixed-order plus parton shower, finding agreement in both the central predictions and their uncertainties at the level expected of the perturbative QCD result. These calculations are in line with the observations of a range of other studies Jäger et al. (2014, 2020); Ballestrero et al. (2018); Höche et al. (2022); Bittrich et al. (2022); Chen et al. (2022a); van Beekveld and Ferrario Ravasio (2024); Barone et al. (2025). In the present paper, we will explore the gluon-gluon fusion background, using a variety of tools in the same common setup as required to obtain theoretically consistent results.
@@
The studies performed in the present work match those used by ATLAS and CMS (referenced earlier) as much as possible. We will show that the nominal setups for the different programs lead to similar predictions for the kinematic distributions crucial to the determination of the VBF Higgs signal. As a result, the uncertainties quoted for the ggF backgrounds by ATLAS and CMS may be overstated. We will compare predictions from approaches that are inclusive in the number of jets (; such as from Sherpa MEPS@LO and MEPS@NLO) to those from more exclusive approaches, where only the final 2 jet state is produced (such as Sherpa S-MC@NLO and Powheg Box + Pythia/ Powheg Box + Herwig ). We also calculate the MiNNLO prediction for gluon-fusion (ggF) Higgs boson production, in which the leading jet spectrum is computed at NLO precision, while the second jet is predicted at LO. This is the framework which is currently used by the ATLAS and CMS experiments. We compare results at the parton shower level and the hadron level and discuss similarities and differences. We disentangle uncertainty estimates as much as possible, in order to provide a better understanding of their sources. We also test the impact of the merging scheme and scale. We provide a Rivet routine that can be directly used by the experiments to confirm the correct use of each generator. Although the ATLAS and CMS determinations of the VBF Higgs cross section are not strictly cut-based, we will also discuss a cutflow analysis, indicating the results of successive phase-space restrictions that are used in VBF Higgs analyses, as well as providing comparisons of what has been termed loose and tight VBF selection cuts.
@@ II Input parameters and Computational Tools
We eliminate the two antipodal combinations and . The parton distribution functions (PDFs) are defined by the central group of PDF4LHC21_40 NNLO PDFs Ball et al. (2022). The computational tools used in this study are introduced in the following. An overview of the perturbative accuracy achieved in each of the tools in the , , and phase spaces is given in Tab. 1. We have cross checked the agreement between NNLOJet, Sherpa, and Powheg Box at NLO. In Fig. 1, we show the rapidity distribution of the Higgs boson as an example.
| Perturbative Accuracy | |||
| Setup | |||
| NNLOJet | - | NNLO | NLO |
| MiNNLO + Pythia/Herwig | NNLO+PS | NLO+PS | LO+PS |
| Sherpa MEPS@NLO | NLO+PS | NLO+PS | NLO+PS |
| Sherpa MC@NLO | - | - | NLO+PS |
| Powheg Box + Pythia/Herwig | - | - | NLO+PS |
@@ II.1 NNLOJet
NNLOJet is a parton level event generator specializing in the computation of NNLO QCD corrections Huss et al. (2025). The antenna subtraction formalism Gehrmann-De Ridder et al. (2005a, b); Daleo et al. (2007, 2010); Gehrmann-De Ridder et al. (2012); Currie et al. (2013) is used to regulate IR divergences at each stage of the fixed-order calculations and to provide fully-differential results. For Higgs production in the gluon-fusion channel, NNLOJet provides predictions at NNLO Chen et al. (2015, 2016). This implies that for observables with at least one (), two () and three jets (), the corresponding fixed order QCD accuracy is NNLO, NLO and LO, respectively. In this study, the calculation is performed in the Higgs effective field theory (HEFT) with infinitely large top-quark mass. The relevant Standard Model (SM) top-quark mass effects were studied for at NLO with full theory approximation on the two-loop matrix elements finding comparable NLO/LO ratios between HEFT and SM for differential observables Chen et al. (2022b), justifying the treatment of top-quark mass in the rest of this study. More details on this can be found in App. C.
@@ II.2 POWHEG BOX
We emphasize that the MiNNLO-based prediction for the 2-jet final state is leading-order accurate. Next-to-leading order precision is achieved by using the simulation, which requires generation cuts or Born suppression (see Sec. IV.2 for details). While Born suppression works well at the parton level, full event generation can fail if the simulation of multiple parton interactions is allowed to generate jets at transverse momenta larger than that of the hard process. Events with low transverse momentum jets in the hard process, which have large (and in principle unbounded) weights, can then be elevated to the 2-jet final state, which will result in distributions with seemingly random statistical outlier events. We therefore decide to define the upper scale for both parton-shower and MPI evolution as the value given by the SCALUP parameter. This technique is not available in Herwig, therefore comparisons including Powheg Box + Herwig will be limited to the parton level.
Note that a sample with all three multiplicities (, and ) provided at NLO accuracy can be obtained with an extension of the MiNLO method described in Ref. Frederix and Hamilton (2016), but the corresponding code is currently not publicly available.
@@ II.3 Pythia
Pythia Sjöstrand et al. (2006); Bierlich et al. (2022) is a multi-purpose particle-level event generator with a historically strong focus on soft physics.‘ The non-perturbative evolution from colored partons to color-singlet hadrons is modeled using its distinctive feature, the string-fragmentation model Andersson et al. (1983); Sjöstrand (1984). In the present version 8.3, Pythia implements the Vincia antenna shower Brooks et al. (2020) in addition to the default -ordered “simple shower” Sjöstrand and Skands (2005)111In versions prior to 8.316, also the Dire dipole-like parton shower Höche and Prestel (2015) was available.. For the simple shower, a dipole-recoil option is available, which replaces the separate incoherent initial-final and final-initial parton evolution by a common, coherent initial-final evolution Cabouat and Sjöstrand (2018), in close correspondence to Vincia. The Apollo parton shower Preuss (2024) is not yet available for initial-state radiation. We therefore limit ourselves to the simple shower with and without using the dipole-recoil option. For each of the parton-shower algorithms, multi-parton interactions are fully interleaved with the shower evolution Sjöstrand and Skands (2005). In this study, we use Pythia 8.315.
Pythia 8.3 implements a series of processes at LO. Both the VBF Higgs-production process and the gluon-fusion Higgs production process in the heavy-top limit are available at LO accuracy. In case of the latter, the simple shower algorithm also incorporates LO matrix-element corrections for the first emission, corresponding to the configuration. For matching and merging, Pythia relies on Les Houches events Boos et al. (2001); Alioli et al. (2014) from external programs such as MadGraph5_aMC@NLO Alwall et al. (2014), Powheg Box, or Sherpa. Multi-jet merging is implemented in the CKKW-L approach Lönnblad (2002); Lönnblad and Prestel (2012). For NLO matching with the Powheg Box, dedicated PowhegHooks are available for all internal parton-shower algorithms to account for the mismatch between the evolution variable in Powheg Box and the shower evolution variable in Pythia Corke and Sjöstrand (2010). NLO matching with MadGraph5_aMC@NLO is only available with the default simple shower with (non-default) global longitudinal recoil in final-state branchings. This was found to be problematic for certain simulations Cabouat and Sjöstrand (2018); Ballestrero et al. (2018); Jäger et al. (2020); Höche et al. (2022) and will therefore not be considered in this study.
@@ II.4 Herwig
In this article, we use the angular-ordered shower to process Les Houches events generated by the Powheg Box. In particular, we use the powhegHerwig plugin developed in Ref. Ferrario Ravasio and Oleari (2023), which is largely based on the Les Houches interface provided by the ThePEG Lonnblad (2009).
@@ II.5 Sherpa
Sherpa Gleisberg et al. (2004, 2009); Bothmann et al. (2019, 2024) is a general-purpose event generator with a strong focus on perturbative physics, in particular multi-jet merging methods. We use version 3.0 of the program for this study. The matching of fixed-order NLO corrections to the parton shower is performed in the S-MC@NLO approach Höche et al. (2012, 2013), while NLO computations are performed using Amegic Krauss et al. (2002); Gleisberg and Krauss (2008), Comix Gleisberg and Höche (2008), and MCFM Campbell and Ellis (1999); Campbell et al. (2011, 2015); Campbell and Neumann (2019); Campbell et al. (2021). The virtual corrections have been cross checked against OpenLoops Cascioli et al. (2012); Buccioni et al. (2018, 2019). We use a modified version of the Catani-Seymour dipole shower algorithm (CSS) Schumann and Krauss (2008), which was cross-checked against the Dire parton shower Höche and Prestel (2015) (for S-MC@NLO only). The Dire shower has been deprecated in Sherpa, while the NLO matching Höche et al. (2025a) for the new Alaric parton shower Herren et al. (2023); Assi and Höche (2024); Höche et al. (2025b) is not yet available for initial-state evolution. Therefore we only present predictions using the default CSS algorithm. The NLO multi-jet merging is based on Gehrmann et al. (2013); Hoeche et al. (2013). Hadronization in Sherpa is provided by a cluster fragmentation model Winter et al. (2004); Chahal and Krauss (2022), and multiple parton scattering as a model of the underlying event is simulated according to Alekhin et al. (2006a, b).
@@ III Comparison to typical LHC simulations
Before performing a detailed analysis of theoretical uncertainties, we compare our new predictions to those obtained from a setup routinely employed by the experimental collaborations. This provides the context for the discussion of the true uncertainties in Sec. IV. The event generator conventionally used by ATLAS and CMS for the ggF Higgs boson production background to VBF Higgs boson production was NNLOPS Hamilton et al. (2013a, 2015). It has now largely been replaced by the more modern MiNNLO generator Monni et al. (2020); Niggetiedt and Wiesemann (2024), which gives formally equivalent results while being substantially more efficient, due to the elimination of the reweighting step. Both NNLOPS and MiNNLO are usually interfaced with the Herwig and Pythia Monte-Carlo programs. The estimated uncertainty in the ggF Higgs boson background is dominated by the observed differences between the predictions of NNLOPS+Pythia and NNLOPS+Herwig, which arise from both perturbative and non-perturbative QCD effects as well as technical differences.
A key point is that these are the differences observed by using the simulations “out of the box” (albeit with tuned non-perturbative parameters from each experiment), and not differences resulting from true discrepancies in the physics models of the event generators. One prominent example of this is the inconsistent usage of the CMW scale scheme Catani et al. (1991), which incorporates the resummed single-logarithmic corrections due to the two-loop cusp anomalous dimension Kodaira and Trentadue (1983) in the parton shower. In Pythia, this correction is optional and it is not included by default, while in Herwig and Sherpa it is enabled by default. The missing single-logarithmic terms in Pythia can be compensated for by using a larger value of the strong coupling, which is often misinterpreted as Pythia actually requiring a different coupling definition. In this section, we will use the ATLAS and CMS setups of MiNNLO for comparison to the ones we recommend as a new baseline to reach the highest theoretical precision in jets from ggF.